If the area of a triangle is 40 cm^2 and the base equals (2h + 6), what is the height?

Here is what I tried.
A=1/2bh
40=1/2(2h + 6)(h)
40= (h^2 + 3h)(h)
40= h^3 + 3h^2
stuck here. Thanks for the help

Answer by stanbon(48535)

If the area of a triangle is 40 cm^2 and the base equals (2h + 6), what is the height?
-----------------------
A=(1/2)bh
---
bh = 2A
h = (2A)/b
--------------------
h = (2*40)/(2h+6)
---
Cross-multiply to get:
2h^2 + 6h - 80 = 0
h^2 +3h - 40 = 0
(h+8)(h-5) = 0
Positive solution:
h = 5 cm
======================
Cheers,
Stan H.